# Revealing Language Model Trajectories via Kullback-Leibler Divergence

### v1

@@ D.2 KL Divergence between Checkpoints
Table 3: KL divergence (bit/byte) between different model sizes at the final checkpoint of Pythia.
Fig. 11 shows the variation in KL divergence between consecutive checkpoints for all model sizes. In this figure, the value at step 10k, for example, represents the KL divergence between the checkpoints at steps 10k and 11k. During pretraining, the learning rate changes dynamically with the training step. In particular, the first 1,430 steps correspond to the warmup phase, during which the learning rate increases linearly. As a result, as shown in Fig. 11, the KL divergence between the checkpoints takes on relatively large values, mostly exceeding 10 bit/byte during this interval. Additionally, Fig. 12 plots the KL divergence between consecutive checkpoints after step 10k for Pythia, along with the corresponding standard error. The derivation of the standard error is described in Appendix C. Finally, Table 3 presents the pairwise KL divergence between model sizes at the final checkpoint. As discussed in Section 3, the divergence caused by differences in model size is larger than that resulting from 1k steps of training.
D.3 Period of Spiral Trajectory
Figure 13: (Left) PC2 vs. PC3 after applying PCA to the later half portion of the 3D t-SNE trajectories for each Pythia model size. (Right) Autocorrelation function of PC3.
In this appendix, we describe the procedure used to analyze the spiral trajectories suggested in Fig. 1. For each model size in Pythia, 3D t-SNE is applied independently. PCA is then performed on the sequence of checkpoints from the midpoint of the trajectory (step 67k onward) to extract principal axes. Since PC1 corresponds to the overall direction of trajectory progression, we plot PC2 versus PC3 in the left panel of Fig. 13, where the trajectory approximately forms a circular pattern. We also compute the autocorrelation function of the PC3 values, treated as a time series, and show the result in the right panel of Fig. 13. The spiral period reported in Table 2 is defined as the step at which the autocorrelation reaches its maximum between the second and third zero-crossings.
D.4 PCA
In Section 3, we visualized the trajectories of Pythia during pretraining using t-SNE. In this appendix, we instead visualize the trajectories using PCA.
Fig. 14 shows the result of applying PCA simultaneously to all model sizes and plotting their trajectories. Fig. 15 further illustrates the evolution of multiple principal component values along the training steps. Unlike t-SNE, the top principal components exhibit similar values across model sizes, resulting in overlapping trajectories. This difference likely arises from the fact that the top principal components capture global variation in the trajectories, whereas t-SNE emphasizes local structure. In contrast, the lower principal components exhibit more noise-like behavior. As shown in Table 3, the KL divergence between different model sizes at the final checkpoint is larger than that between consecutive checkpoints. This is consistent with the observation that trajectories are separated in the lower components.
D.5 Comparison with Weights
Figure 16: t-SNE visualization of the trajectories based on weight-space distances for Pythia models after step 100k. To align scales across model sizes, the trajectories were centered after t-SNE. They were also rotated to align the directions of progression. Final checkpoints are indicated using distinct markers.
@@ D.5 Comparison with Weights
This appendix describes the procedure for comparing language models using the Euclidean distance between the weights of Pythia checkpoints.
Trajectory.
For each model size, we computed a distance matrix containing the pairwise Euclidean distances among all checkpoints. Fig. 16 shows the trajectories from step 100k onward, visualized using t-SNE999t-SNE can take a distance matrix as input.. The trajectories were centered after t-SNE to align the plot ranges. In addition, for the 1B model, the trajectory was rotated by 180 degrees to align its rough direction of progression with that of the other model sizes.
Anomalous diffusion.
Fig. 3 shows the evolution of squared Euclidean distances in weights and log-likelihood vectors over 10k steps starting from step 120k. Fig. 4 illustrates how the estimated diffusion exponent changes when varying the starting step for the calculation. To ensure stability, we vary the starting point only after the warmup phase. The exponent is estimated using linear regression (least squares) over a 10k-step window beginning at each initial step. Summary statistics of the coefficient of determination ($R^{2}$) for each regression are provided in Table 4.
D.6 Different Seeds
@@
van der Wal et al. (2025) trained Pythia models with nine different seeds and released checkpoints at the same steps as in the original Pythia suite. The random seed affects the initialization of weights and the order of the training data. We used all available checkpoints for the 410M models with seeds 1 through 9. However, at the time of writing this paper, the weights for (seed, step) = (2, 111k), (6, 88k), and (9, 40k) were not available on Hugging Face and were therefore excluded from our analysis [^10] For example, see https://huggingface.co/EleutherAI/pythia-410m-seed2/tree/step111000.. The original 410M model was trained with seed 1234.
@@
Fig. 17 shows the KL divergence between the final checkpoints of models trained with different seeds. Despite sharing the same architecture and data, the KL divergence between seed 3 and the other seeds, as well as between seed 4 and the others, are significantly larger. According to van der Wal et al. (2025), models trained with seed 3 and seed 4 experienced loss spikes during training and are considered outliers in terms of performance. This suggests that KL divergence is also effective in identifying such outlier models. Therefore, seed 3 and seed 4 are excluded when generating Fig. 1. We also report in Table 1(a) the average KL divergence between the final checkpoint of the original 410M model and those of seeds 1, 2, 5, 6, 7, 8, and 9.
Appendix E Details of Section 4
Figure 18: Cumulative distribution of KL divergence for pairs of fine-tuned models and their base models, and for randomly selected model pairs (corresponding to Table 1(b)). The horizontal axis is scaled as $\log_{10}(1+\mathrm{KL})$. The dotted lines indicate the median of each distribution.
We describe the language models used in Section 4. From the 1,018 models analyzed in Oyama et al. (2025), we identified fine-tuned models as those whose base model is specified in the metadata available via the Hugging Face Hub API [^11] https://github.com/huggingface/huggingface_hub. Our analysis was conducted on pairs of fine-tuned models and their base models. Models associated with multiple base models (i.e., merged models) were excluded. A list of the language models used in this analysis is provided in Table LABEL:tab:model_list_ft in Appendix G. In Fig. 5, there are a few pairs connected by line segments that span across different colors, namely different model types. This occurs because, under the model type classification method used in Oyama et al. (2025), some fine-tuned models were categorized as belonging to the “others” type. In our experiments, when we refer to, for example, a “llama-3 fine-tuned pair”, we mean a pair in which the fine-tuned model is of the llama-3 type.
Fig. 18 illustrates the cumulative distributions of KL divergence for pairs of fine-tuned models and their base models, as well as for randomly selected model pairs. Summary statistics for these distributions are reported in Table 1(b). The figure shows that KL divergence after fine-tuning tends to be slightly smaller, as indicated by the left-skewed distribution. As noted in Oyama et al. (2025), models of the same type tend to have smaller KL divergence. To more accurately assess the effect of fine-tuning, we also analyze random model pairs sampled within the same model type.
Appendix F Details of Section 5
Figure 19: Re-rendering of the top panel of Fig. 6. Colors indicate mean log-likelihood, clipped at the bottom 5% across all values to improve visibility.
Figure 20: 3D t-SNE visualization of the left panel of Fig. 7. Deeper layers are represented with lighter colors, and the final layer is indicated by a white square.
@@ Appendix F Details of Section 5
Here, we describe the language models used in Section 5. For the trajectories shown in Fig. 6 and Fig. 19, we selected the 50 most downloaded models from those analyzed in Oyama et al. (2025). For specified model type analysis, we selected models with 32 layers from the llama-2, llama-3, and mistral types. For each of these types, we used the 10 most downloaded models. These models were used in the analysis shown in Fig. 7 and Fig. 20. A complete list of the models is provided in Table LABEL:tab:model_list_layer in Appendix G.
@@
Fig. 19 shows the trajectories of models across layers, color-coded by mean log-likelihood, depending on layer depth. The mean log-likelihood increases as the layer depth increases. This indicates that the direction of trajectory progression corresponds to deeper layers, that is, to increasing mean log-likelihood. Additionally, Fig. 20 presents a 3D version of the left panel of Fig. 6. This illustration implies that the jump observed in Fig. 7 is not an artifact of the 2D t-SNE visualization, but rather results from a genuinely large KL divergence between adjacent layers.
Appendix G Model List
@@
The models used in Sections 3, 4, and 5 are listed in Tables LABEL:tab:model_list_pretraining, LABEL:tab:model_list_ft, and LABEL:tab:model_list_layer, respectively. Table LABEL:tab:model_list_pretraining is sorted by model size; Table LABEL:tab:model_list_ft is sorted alphabetically by model name, considering the model generation, model type, and corresponding base model; and Table LABEL:tab:model_list_layer is sorted in descending order of download counts.
@@
BibTeX entries for each model were selected according to the procedure described by Oyama et al. (2025). Note that, for the Pythia models trained with different random seeds van der Wal et al. (2025) and used in Section 3, the corresponding BibTeX entries in Table LABEL:tab:model_list_pretraining were prepared manually.
@@
Table 5: List of 14 models in the experiments for pretraining models. “ID” denotes the index sorted by model size; “Model Name” denotes the name of the model.
@@
1 EleutherAI/pythia-410m Gao et al. (2020); Biderman et al. (2022, 2023)
2 EleutherAI/pythia-410m-seed1 van der Wal et al. (2025)
3 EleutherAI/pythia-410m-seed2 van der Wal et al. (2025)
4 EleutherAI/pythia-410m-seed3 van der Wal et al. (2025)
5 EleutherAI/pythia-410m-seed4 van der Wal et al. (2025)
6 EleutherAI/pythia-410m-seed5 van der Wal et al. (2025)
7 EleutherAI/pythia-410m-seed6 van der Wal et al. (2025)
8 EleutherAI/pythia-410m-seed7 van der Wal et al. (2025)
9 EleutherAI/pythia-410m-seed8 van der Wal et al. (2025)
10 EleutherAI/pythia-410m-seed9 van der Wal et al. (2025)
11 EleutherAI/pythia-1b Gao et al. (2020); Biderman et al. (2022, 2023)
12 EleutherAI/pythia-1.4b Gao et al. (2020); Biderman et al. (2022, 2023)
13 EleutherAI/pythia-2.8b Gao et al. (2020); Biderman et al. (2022, 2023)
14 EleutherAI/pythia-6.9b Gao et al. (2020); Biderman et al. (2022, 2023)
@@
1 deepseek-ai/DeepSeek-Prover-V1.5-Base Xin et al. (2024) deepseek 1 230
2 tiiuae/falcon-rw-1b Brown et al. (2020); Press et al. (2022); Dao et al. (2022); Penedo et al. (2023) falcon 1 22921
3 google/gemma-1.1-7b-it Joshi et al. (2017); Mihaylov et al. (2018); Zhao et al. (2018); Clark et al. (2019); Talmor et al. (2019); Sap et al. (2019); Sakaguchi et al. (2019); Bisk et al. (2019); Chollet (2019); Zellers et al. (2019); Hendrycks et al. (2021); Chen et al. (2021); Austin et al. (2021); Cobbe et al. (2021); Parrish et al. (2022); Srivastava et al. (2023); Zhong et al. (2023); Gemini Team et al. (2024) gemma 1 19835
4 google/gemma-7b Joshi et al. (2017); Mihaylov et al. (2018); Zhao et al. (2018); Rudinger et al. (2018); Bisk et al. (2019); Chollet (2019); Clark et al. (2019); Talmor et al. (2019); Sap et al. (2019); Zellers et al. (2019); Sakaguchi et al. (2019); Gehman et al. (2020); Chen et al. (2021); Austin et al. (2021); Hendrycks et al. (2021); Dhamala et al. (2021); Cobbe et al. (2021); Lin et al. (2022); Hartvigsen et al. (2022); Parrish et al. (2022); Srivastava et al. (2023); Zhong et al. (2023); Dettmers et al. (2023); Gemini Team et al. (2024) gemma 1 72047
5 openchat/openchat-3.5-0106-gemma Wang et al. (2024a) gemma 1 7817
6 openlm-research/open_llama_3b Together Computer (2023); Touvron et al. (2023a); Geng and Liu (2023) llama-1 1 157405
7 codellama/CodeLlama-13b-Instruct-hf Rozière et al. (2024) llama-2 1 25248
8 meta-llama/Llama-2-13b-hf Touvron et al. (2023b) llama-2 1 120871
9 meta-llama/Llama-2-7b-hf Touvron et al. (2023b) llama-2 1 1294737
10 meta-llama/Meta-Llama-3-8B-Instruct AI@Meta (2024) llama-3 1 2058011
11 meta-llama/Meta-Llama-3-8B AI@Meta (2024) llama-3 1 675850
12 NousResearch/Hermes-2-Pro-Llama-3-8B "Teknium" et al. (2024b) llama-3 1 26797
13 mistralai/Mistral-7B-Instruct-v0.2 Jiang et al. (2023) mistral 1 3565248
14 mistralai/Mistral-7B-v0.1 Jiang et al. (2023) mistral 1 1750089
@@
17 stabilityai/japanese-stablelm-base-gamma-7b Jiang et al. (2023) mistral 1 2072
@@
19 microsoft/Orca-2-13b Mitra et al. (2023) others 1 13616
20 upstage/SOLAR-10.7B-v1.0 Kim et al. (2024b) others 1 24478
21 01-ai/Yi-1.5-9B 01. AI et al. (2025) others 1 20252
22 deepseek-ai/DeepSeek-Prover-V1.5-SFT Xin et al. (2024) deepseek-ai/DeepSeek-Prover-V1.5-Base deepseek 2 6459
@@
26 google/gemma-7b-it Joshi et al. (2017); Mihaylov et al. (2018); Zhao et al. (2018); Rudinger et al. (2018); Bisk et al. (2019); Chollet (2019); Clark et al. (2019); Talmor et al. (2019); Sap et al. (2019); Zellers et al. (2019); Sakaguchi et al. (2019); Gehman et al. (2020); Chen et al. (2021); Austin et al. (2021); Hendrycks et al. (2021); Dhamala et al. (2021); Cobbe et al. (2021); Lin et al. (2022); Hartvigsen et al. (2022); Parrish et al. (2022); Srivastava et al. (2023); Zhong et al. (2023); Gemini Team et al. (2024) google/gemma-7b gemma 2 66776
@@
29 rinna/youri-7b Andonian et al. (2021); Touvron et al. (2023b); Zhao et al. (2023); Sawada et al. (2024) meta-llama/Llama-2-7b-hf llama-2 2 2512
30 aaditya/Llama3-OpenBioLLM-8B Singhal et al. (2022); Nori et al. (2023); Singhal et al. (2023); Pal and Sankarasubbu (2024b, a); Rafailov et al. (2024) meta-llama/Meta-Llama-3-8B llama-3 2 9308

### v2

@@ E.2 KL Divergence between Checkpoints
Table 2: KL divergence (bits/byte) between different model sizes at the final checkpoint of Pythia.
During pretraining, the learning rate changes dynamically with the training step. In particular, the first 1,430 steps correspond to the warmup phase, during which the learning rate increases linearly. As a result, as shown in the left panel of Fig. 4, the KL divergence between the checkpoints takes on relatively large values, mostly exceeding 10 bits/byte during this interval. The right panel of Fig. 4 shows the variation in KL divergence between consecutive saved checkpoints for all model sizes. In this figure, the value at step 10k, for example, represents the KL divergence between the checkpoints at steps 10k and 11k. Additionally, Fig. 14 plots the KL divergence between consecutive saved checkpoints after step 10k for Pythia, along with the corresponding standard error. The derivation of the standard error is described in Appendix C. Finally, Table 2 presents the pairwise KL divergence between model sizes at the final checkpoint. As discussed in Section 4, the divergence caused by differences in model size is larger than that resulting from 1k steps of training.
E.3 Estimation of Diffusion Exponents
@@
In this subsection, we describe the procedure for estimating diffusion exponents in both the log-likelihood and weight spaces. The left panel of Fig. 5 shows the evolution of squared Euclidean distances in weights and log-likelihood vectors over 10k steps starting from step 120k. The right panel of Fig. 5 illustrates how the estimated diffusion exponent changes when varying the starting step for the calculation. To ensure stability, we vary the starting point only after the warmup phase. The exponent is estimated using linear regression (least squares) over a 10k-step window beginning at each initial step. Summary statistics of the coefficient of determination ($R^{2}$) for each regression are provided in Table 3.
E.4 Different Seeds
@@
wal2025polypythias trained Pythia models with nine different seeds and released checkpoints at the same steps as in the original Pythia suite. The random seed affects the initialization of weights and the order of the training data. We used all available checkpoints for the 410M models with seeds 1 through 9. However, at the time of writing this paper, the weights for (seed, step) = (2, 111k), (6, 88k), and (9, 40k) were not available on Hugging Face and were therefore excluded from our analysis [^11] For example, see https://huggingface.co/EleutherAI/pythia-410m-seed2/tree/step111000.. The original 410M model was trained with seed 1234.
@@
Fig. 15 shows the KL divergence between the final checkpoints of models trained with different seeds. Despite sharing the same architecture and data, the KL divergence between seed 3 and the other seeds, as well as between seed 4 and the others, are significantly larger. According to wal2025polypythias, models trained with seed 3 and seed 4 experienced loss spikes during training and are considered outliers in terms of performance. This suggests that KL divergence is also effective in identifying such outlier models. Therefore, seed 3 and seed 4 are excluded when generating Fig. 1. We also report in Table 1(b) the average KL divergence between the final checkpoint of the original 410M model and those of seeds 1, 2, 5, 6, 7, 8, and 9.
Appendix F Details of Section 3.2
@@
Here, we describe the language models used in Section 3.2. We selected the 50 most downloaded models from those analyzed in Oyama et al. (2025). A complete list of the models is provided in Table LABEL:tab:model_list_layer in Appendix I.
For the quantization methods, we applied 8-bit and 4-bit quantization using bitsandbytes [^12] https://github.com/bitsandbytes-foundation/bitsandbytes Dettmers et al. (2022).
Appendix G Details of Section 3.3
We describe the language models used in Section 3.3. From the 1,018 models analyzed in Oyama et al. (2025), we identified fine-tuned models as those whose base model is specified in the metadata available via the Hugging Face Hub API [^13] https://github.com/huggingface/huggingface_hub. Our analysis was conducted on pairs of fine-tuned models and their base models. Models associated with multiple base models (i.e., merged models) were excluded. A list of the language models used in this analysis is provided in Table LABEL:tab:model_list_ft in Appendix I. In Fig. 2b, there are a few pairs connected by line segments that span across different colors, namely different model types. This occurs because, under the model type classification method used in Oyama et al. (2025), some fine-tuned models were categorized as belonging to the “others” type.
Summary statistics for these distributions are reported in Table 1(a). Fig. 3 shows that KL divergence after fine-tuning tends to be slightly smaller, as indicated by the left-skewed distribution. As noted in Oyama et al. (2025), models of the same type tend to have smaller KL divergence. To more accurately assess the effect of fine-tuning, we also analyze random model pairs sampled within the same model type.
Appendix H Details of Section 3.4
Figure 16: Re-rendering of the top panel of Fig. 2c. Colors indicate mean log-likelihood, clipped at the bottom 5% across all values to improve visibility.
Figure 17: For 10 models per model type: (left) t-SNE visualization of trajectories across layers, where the final layer is indicated by a white square; (right) KL divergence between adjacent layers (corresponding to Table 1(c)).
Figure 18: 3D t-SNE visualization (perplexity=30) of the left panel of Fig. 17. Deeper layers are represented with lighter colors, and the final layer is indicated by a white square.
Here, we describe the language models used in Section 3.4. For the trajectories shown in Fig. 2c and Fig. 16, we selected the 50 most downloaded models from those analyzed in Oyama et al. (2025). A complete list of the models is provided in Table LABEL:tab:model_list_layer in Appendix I.
@@
Fig. 16 shows the trajectories of models across layers, color-coded by mean log-likelihood, depending on layer depth. The mean log-likelihood increases as the layer depth increases. This indicates that the direction of trajectory progression corresponds to deeper layers, that is, to increasing mean log-likelihood.
Additionally, for detailed analysis, we selected models with 32 layers from the llama-2, llama-3, and mistral types. For each of these types, we used the 10 most downloaded models. That is, the top 50 do not include 10 models for each of llama-2, llama-3, and mistral types. Therefore, additional models were included to ensure 10 models per category, resulting in a total of 9 added models. Fig. 17 shows the trajectories of these models and the KL divergence between adjacent layers, while Fig. 18 illustrates the trajectories in 3d space. For all model types, the trajectories are already clustered by model type from the shallow layers onward. In addition, the KL divergence values exhibit similar trends within each type.
Appendix I Model List
@@
The models used in Sections 3.1, 3.3, and 3.4 are listed in Tables LABEL:tab:model_list_pretraining, LABEL:tab:model_list_ft, and LABEL:tab:model_list_layer, respectively. Table LABEL:tab:model_list_pretraining is sorted by model size; Table LABEL:tab:model_list_ft is sorted alphabetically by model name, considering the model generation, model type, and corresponding base model; and Table LABEL:tab:model_list_layer is sorted in descending order of download counts.
@@
BibTeX entries for each model were selected according to the procedure described by Oyama et al. (2025). Note that, for the Pythia models trained with different random seeds wal2025polypythias and used in Section 4, the corresponding BibTeX entries in Table LABEL:tab:model_list_pretraining were prepared manually.
@@
Table 4: List of 14 models in the experiments for pretraining models. “ID” denotes the index sorted by model size; “Model Name” denotes the name of the model. Seeds 3 and 4 of Pythia-410M are excluded from the experiments in Sections 3.1 and 4.
@@
1 EleutherAI/pythia-410m arxiv:2101.00027; arxiv:2201.07311; arxiv:2304.01373
2 EleutherAI/pythia-410m-seed1 wal2025polypythias
3 EleutherAI/pythia-410m-seed2 wal2025polypythias
4 EleutherAI/pythia-410m-seed3 wal2025polypythias
5 EleutherAI/pythia-410m-seed4 wal2025polypythias
6 EleutherAI/pythia-410m-seed5 wal2025polypythias
7 EleutherAI/pythia-410m-seed6 wal2025polypythias
8 EleutherAI/pythia-410m-seed7 wal2025polypythias
9 EleutherAI/pythia-410m-seed8 wal2025polypythias
10 EleutherAI/pythia-410m-seed9 wal2025polypythias
11 EleutherAI/pythia-1b arxiv:2101.00027; arxiv:2201.07311; arxiv:2304.01373
12 EleutherAI/pythia-1.4b arxiv:2101.00027; arxiv:2201.07311; arxiv:2304.01373
13 EleutherAI/pythia-2.8b arxiv:2101.00027; arxiv:2201.07311; arxiv:2304.01373
14 EleutherAI/pythia-6.9b arxiv:2101.00027; arxiv:2201.07311; arxiv:2304.01373
@@ 10 EleutherAI/pythia-410m-seed9 wal2025polypythias
1 deepseek-ai/DeepSeek-Prover-V1.5-Base arxiv:2408.08152 deepseek 1 230
2 tiiuae/falcon-rw-1b arxiv:2005.14165; arxiv:2108.12409; arxiv:2205.14135; arxiv:2306.01116 falcon 1 22921
3 google/gemma-1.1-7b-it arxiv:1705.03551; arxiv:1809.02789; arxiv:1804.06876; arxiv:1905.10044; arxiv:1811.00937; arxiv:1904.09728; arxiv:1907.10641; arxiv:1911.11641; arxiv:1911.01547; arxiv:1905.07830; arxiv:2009.03300; arxiv:2107.03374; arxiv:2108.07732; arxiv:2110.14168; arxiv:2110.08193; arxiv:2206.04615; arxiv:2304.06364; arxiv:2312.11805 gemma 1 19835
4 google/gemma-7b arxiv:1705.03551; arxiv:1809.02789; arxiv:1804.06876; arxiv:1804.09301; arxiv:1911.11641; arxiv:1911.01547; arxiv:1905.10044; arxiv:1811.00937; arxiv:1904.09728; arxiv:1905.07830; arxiv:1907.10641; arxiv:2009.11462; arxiv:2107.03374; arxiv:2108.07732; arxiv:2009.03300; arxiv:2101.11718; arxiv:2110.14168; arxiv:2109.07958; arxiv:2203.09509; arxiv:2110.08193; arxiv:2206.04615; arxiv:2304.06364; arxiv:2305.14314; arxiv:2312.11805 gemma 1 72047
5 openchat/openchat-3.5-0106-gemma arxiv:2309.11235 gemma 1 7817
6 openlm-research/open_llama_3b together2023redpajama; arxiv:2302.13971; openlm2023openllama llama-1 1 157405
7 codellama/CodeLlama-13b-Instruct-hf arxiv:2308.12950 llama-2 1 25248
8 meta-llama/Llama-2-13b-hf arxiv:2307.09288 llama-2 1 120871
9 meta-llama/Llama-2-7b-hf arxiv:2307.09288 llama-2 1 1294737
10 meta-llama/Meta-Llama-3-8B-Instruct llama3modelcard llama-3 1 2058011
11 meta-llama/Meta-Llama-3-8B llama3modelcard llama-3 1 675850
12 NousResearch/Hermes-2-Pro-Llama-3-8B Hermes-2-Pro-Llama-3-8B llama-3 1 26797
13 mistralai/Mistral-7B-Instruct-v0.2 arxiv:2310.06825 mistral 1 3565248
14 mistralai/Mistral-7B-v0.1 arxiv:2310.06825 mistral 1 1750089
@@ 12 NousResearch/Hermes-2-Pro-Llama-3-8B Hermes-2-Pro-Llama-3-8B llama-3 1 26797
17 stabilityai/japanese-stablelm-base-gamma-7b arxiv:2310.06825 mistral 1 2072
@@
19 microsoft/Orca-2-13b arxiv:2311.11045 others 1 13616
20 upstage/SOLAR-10.7B-v1.0 arxiv:2312.15166 others 1 24478
21 01-ai/Yi-1.5-9B arxiv:2403.04652 others 1 20252
22 deepseek-ai/DeepSeek-Prover-V1.5-SFT arxiv:2408.08152 deepseek-ai/DeepSeek-Prover-V1.5-Base deepseek 2 6459
@@
26 google/gemma-7b-it arxiv:1705.03551; arxiv:1809.02789; arxiv:1804.06876; arxiv:1804.09301; arxiv:1911.11641; arxiv:1911.01547; arxiv:1905.10044; arxiv:1811.00937; arxiv:1904.09728; arxiv:1905.07830; arxiv:1907.10641; arxiv:2009.11462; arxiv:2107.03374; arxiv:2108.07732; arxiv:2009.03300; arxiv:2101.11718; arxiv:2110.14168; arxiv:2109.07958; arxiv:2203.09509; arxiv:2110.08193; arxiv:2206.04615; arxiv:2304.06364; arxiv:2312.11805 google/gemma-7b gemma 2 66776
@@
29 rinna/youri-7b gpt-neox-library; arxiv:2307.09288; rinna-youri-7b; arxiv:2404.01657 meta-llama/Llama-2-7b-hf llama-2 2 2512
30 aaditya/Llama3-OpenBioLLM-8B arxiv:2212.13138; arxiv:2303.13375; arxiv:2305.09617; OpenBioLLMs; arxiv:2402.07023; arxiv:2305.18290 meta-llama/Meta-Llama-3-8B llama-3 2 9308